A football quarterback is moving straight backward at a speed of 2.00 m/s when he throws a pass to a player 18.0 m straight downfield. The ball is thrown at an angle of 25.0
Answer #1

:  Vox =- Vo cos 25
Voy = V0 sin 25

X = {V0^2sin 2(25)}/g = V0^2*0.77/9.8
V0^2 = sqrt(229). V0 = 15.1 m/s This is the initial speed.